The Insurance Institute for Highway Safety publishes data on the total damage suffered by compact automobiles in a series of controlled, low-speed collisions. The cost for a sample of nine cars, in hundreds of dollars, is provided below[10 6 8 10 4 3.5 7.5 8 9] What is the median cost of the total damage suffered for this sample of cars?

Question 18 options:

$730

$1000

$400

$800
Answer #1

Answer is $400

●Meadian = (n + 1) / 2 = ( 9 + 1 )/2 = 5th Observation

So 5th Observation is 4.